### v1

# Cross-Domain Graph Data Scaling: A Showcase with Diffusion Models

Abstract
Models for natural language and images benefit from data scaling behavior: the more data fed into the model, the better they perform. This ’better with more’ phenomenon enables the effectiveness of large-scale pre-training on vast amounts of data. However, current graph pre-training methods struggle to scale up data due to heterogeneity across graphs. To achieve effective data scaling, we aim to develop a general model that is able to capture diverse data patterns of graphs and can be utilized to adaptively help the downstream tasks. To this end, we propose UniAug, a universal graph structure augmentor built on a diffusion model. We first pre-train a discrete diffusion model on thousands of graphs across domains to learn the graph structural patterns. In the downstream phase, we provide adaptive enhancement by conducting graph structure augmentation with the help of the pre-trained diffusion model via guided generation. By leveraging the pre-trained diffusion model for structure augmentation, we consistently achieve performance improvements across various downstream tasks in a plug-and-play manner. To the best of our knowledge, this study represents the first demonstration of a data-scaling graph structure augmentor on graphs across domains. Our implementations are publicly available on https://github.com/WenzhuoTang/UniAug.
1 Introduction
The effectiveness of existing foundation models radford2021learning ; touvron2023llama ; kirillov2023segment heavily relies on the availability of substantial amounts of data, where the relationship manifests as a scaling behavior between model performance and data scale kaplan2020scaling . Consistent performance gain has been observed with the increasing scale of pre-training data in both Natural Language Processing kaplan2020scaling ; hoffmann2022training and Computer Vision abnar2022exploring ; zhai2022scaling domains. This data scaling phenomenon facilitates the development of general models endowed with extensive knowledge and effective data pattern recognition capabilities. In downstream applications, these models are capable of adaptively achieving performance gains across different tasks.
In the context of graphs, the availability of large-scale graph databases rossi2015network ; hu2020OGB ; snapnets enables possible data scaling across datasets and domains. Existing works have demonstrated graph data scaling following two limited settings: in-domain pre-training xia2023molebert ; liu2024rethinking and task-specific selection for pre-training data cao2023pre . During the pre-training process, each graph in the pre-training pool must be validated as in-domain or relevant to the downstream dataset. Given a specific domain or task, the crucial discriminative data patterns are likely confined to a fixed set mao2024revisiting , leaving other potential patterns in diverse graph data distribution as noisy input. In terms of structure, graphs from different domains are particularly composed of distinct patterns milo2002network , making it hard to transfer across domains. For example, considering the building blocks of the graphs, the motifs shared by ecological food webs differ significantly from those shared by genetic networks milo2002network . Therefore, closely aligning the characteristics of the pre-training graphs and the downstream data both in feature and structure is essential for facilitating positive transfer cao2023pre . As a consequence, the necessity of such meticulous data filtering restricts these methods to scaling up graphs effectively, as they can only utilize a small part of the available data. Given the limitation of the graph pre-training methods, a pertinent question emerges: How can we effectively leverage the increasing scale of graph data across domains?
Rather than focusing solely on data patterns specific to particular domains, we aim to develop a model that has a comprehensive understanding of data patterns inherent across various types of graphs. In line with the principles of data scaling, we hypothesize that incorporating a broader range of training datasets can help the model build an effective and universal graph pattern library, avoiding an overemphasis on major data patterns specific to any single dataset mao2024demystifying . To construct such a general-purpose model, we propose to utilize a diffusion model operating only on the structure as the backbone, for the following key reasons. (1) Current supervised and self-supervised methods tend to capture only specific patterns of graph data, with models designed for particular inductive biases mao2024demystifying ; mao2024revisiting ; xu2018powerful . For instance, graph convolutional networks (GCNs) excel in node-level representation learning by emphasizing homophily, whereas graph-level representation learning benefits from expressive GNNs capable of distinguishing complex graph structures. (2) We opt for a structure-only model due to the heterogeneous feature spaces across graphs, which often include missing features or mismatched semantics mao2024graph . For instance, node features yield completely different interpretations in citation networks, where they represent keywords of documents, compared to molecular networks, where they denote properties of atoms. To this end, we pre-train a structure-only diffusion model on thousands of graphs, which serves as the upstream component of our framework.
In the downstream stage, we employ the pre-trained diffusion model as a Universal graph structure Augmentor (UniAug) to enhance the dataset, where diffusion guidance ho2022classifier ; dhariwal2021diffusion ; gruver2024protein is employed to align the generated structure with the downstream requirements. Specifically, we generate synthetic structures with various guidance objectives, and the resulting graphs consist of generated structures and original node features. This data augmentation paradigm strategically circumvents feature heterogeneity and fully utilizes downstream inductive biases by applying carefully designed downstream models to the augmented graphs in a plug-and-play manner. Empirically, we apply UniAug to graphs from diverse domains and consistently observe performance improvement in node classification, link prediction, and graph property prediction. To the best of our knowledge, this study represents the first demonstration of a data-scaling graph structure augmentor on graphs across domains.
2 Preliminary and Related Work
Learning from unlabeled graphs Graph self-supervised learning (SSL) methods provide examples of pre-training and fine-tuning paradigm hu2019strategies ; hou2022graphmae ; kim2022graph ; you2021graph ; xu2021self . However, these methods benefit from limited data scaling due to feature heterogeneity, structural pattern differences across domains, and varying downstream inductive biases. It is worth mentioning that DCT liu2024data presents a pre-training and then data augmentation pipeline on molecules. Despite its impressive performance improvement on graph-level tasks, DCT is bounded with molecules and thus the use cases are limited.
Graph data augmentation There have been many published works exploring graph data augmentation (GDA) since the introduction of graph neural networks (GNNs), with a focus on node-level park2021metropolis ; liu2022local ; azabou2023half , link-level zhao2022learning ; nguyen2024diffusion , and graph-level han2022g ; ling2023graph ; luo2022automated ; liu2022graph ; kong2022robust . These GDA methods have been generally designed for specific tasks or particular aspects of graph data. In addition, they are often tailored for a single dataset and struggle to transfer to unseen patterns, which limits their generalizability to a broader class of applications.
Diffusion models on graphs Diffusion models ho2020denoising ; song2021scorebased ; rombach2022high are latent variable models that learn data distribution by gradually adding noise into the data and then recovering the clean input. Existing diffusion models on graphs can be classified into two main categories depending on the type of noise injected, i.e. Gaussian or discrete. Previous works employed Gaussian diffusion models both on general graphs niu2020permutation ; jo2022score and molecules shi2021learning ; xu2022geodiff . However, adding Gaussian noise into the adjacency matrix will destroy the sparsity of the graph, which hinders the scalability of the diffusion models haefeli2022diffusion . Recent works adapted discrete diffusion models to graphs with categorical transition kernels vignac2023digress ; chen2023efficient ; chen2023dexplainer . We denote the adjacency matrix of a graph as $\mathbf{A}^{0}\in\{0,1\}^{n\times n}$ with $n$ nodes. With details in Appendix A, we write the forward process to corrupt the adjacency matrix into a sequence of latent variables as Bernoulli distribution
[…]
where $\pi$ is the converging non-zero probability, $\alpha^{t}$ is the noise scale, and $\bar{\alpha}^{t}=\prod_{i=1}^{t}\alpha^{i}$. Under predict-$\mathbf{A}^{0}$ parameterization, the reverse process denoise the adjacency matrix with a Markov chain
$p_{\theta}\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t}\right)\propto\sum_{\widetilde{\mathbf{A}}_{0}}q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t},\widetilde{\mathbf{A}}_{0}\right)\tilde{p}_{\theta}\left(\widetilde{\mathbf{A}}_{0}\mid\mathbf{A}^{t}\right),$
where $\tilde{p}_{\theta}(\widetilde{\mathbf{A}}_{0}\mid\mathbf{A}^{t})$ represents the denoising network that predicts the original adjacency matrix from the noisy adjacency matrix. The parameters are estimated by optimizing the variational lower bound on the negative log-likelihood austin2021structured
[…]
3 Method
Figure 1: The pipeline of UniAug. We pre-train a diffusion model across domains and perform structure augmentation on the downstream graphs. The augmented graphs consist of generated structures and original node features and are then processed by a downstream GNN.
In this section, our goal is to build UniAug to understand the diverse structure patterns of graphs and perform data augmentation with a range of objectives. As illustrated in Fig.1, UniAug consists of two main components: a pre-trained diffusion model and the downstream adaptation through structure augmentation. We first collect thousands of graphs from varied domains with diverse patterns. To construct a general model free of downstream inductive biases, we train a self-conditioned discrete diffusion model on graph structures. In the downstream stage, we train an MLP guidance head on top of the diffusion model with objectives across different levels of granularity. We then augment the downstream dataset by generating synthetic structures through guided generation, where the augmented graph is composed of generated structures and original node features. Subsequently, we apply the augmented data to train a task-specific model for performing downstream tasks. Below, we elaborate on the data collection process, the architecture of the discrete diffusion model, and the guidance objectives employed.
3.1 Pre-training data collection
Figure 2: Normalized structural properties.
In light of the data scaling spirit, we expect our pre-training data to contain diverse data patterns with sufficient volume. As graphs from different domains exhibit different patterns milo2002network , we wish to build a collection of graphs from numerous domains to enable a universal graph pattern library with pre-training. Within the publicly available graph databases, Network Repository rossi2015network provides a comprehensive collection of graphs with varied scales from different domains, such as biological networks, chemical networks, social networks, and many more. Among the thousands of graphs in the Network Repository, some of them contain irregular patterns, including multiple levels of edges, extremely high density, et cetera. To ensure the quality of the graphs, we analyze the graph properties following Xu et al. xu2023better and filter out the outliers. In addition, we observe that the coverage of graphs in the Network Repository is incomplete according to the network entropy and scale-free exponent, as we observe a relatively scattered space in the middle of Fig. 2. To fill in the gap, we include a $1000$ graphs subset of the GitHub Star dataset rozemberczki2020karate . The selected graphs are utilized to train a discrete diffusion model.
3.2 Pre-training through diffusion model
Diffusion models have demonstrated the ability to facilitate transferability from a data augmentation perspective on the images trabucco2024effective ; you2024diffusion ; he2023is . Unlike the traditional hand-crafted data augmentation methods, diffusion models are capable of producing more diverse patterns with high quality trabucco2024effective . With the aid of diffusion guidance ho2022classifier ; dhariwal2021diffusion , these methods can achieve domain customization tailored to specific semantic spaces you2024diffusion ; he2023is . Despite the success of data augmentation through diffusion models on images, the non-Euclidean nature of graph structures poses challenges for data-centric learning on graphs. In addition, the fact that most graphs in the Network Repository are unlabeled exacerbates the challenges, as the absence of labeled data results in substantially lower generation quality for diffusion models dhariwal2021diffusion ; bao2022conditional .
To address the aforementioned challenges, we propose to construct a self-conditioned discrete diffusion model on graph structures. Unlike Gaussian-based diffusion models, discrete diffusion models hoogeboom2021argmax ; austin2021structured ; campbell2022continuous ; vignac2023digress operate with discrete transition kernels between latent variables, as shown in Section 2. The key reason we opt for the discrete diffusion models lies in the sparse nature of graphs, where adding Gaussian noise into the adjacency matrix will result in a dense graph haefeli2022diffusion . On the contrary, discrete diffusion models effectively preserve the sparse structure of graphs during the diffusion process, thus maintaining the efficiency of the models on graphs.
To accommodate for unlabeled graphs, we adopt a self-supervised labeling strategy as an auxiliary conditioning procedure gao2022large ; hu2023self . By leveraging the self-labeling technique, we are able to upscale the diffusion model to data with more diverse patterns gao2022large . The self-labeling technique requires two components, i.e., a feature extractor and a self-supervised annotator.
Feature extractor. We extract graph-level features by calculating graph properties, including the number of nodes, density, network entropy, average degree, degree variance, and scale-free exponent following Xu et al. xu2023better . The first two represent the scale of the graph corresponding to nodes and edges, and the rest indicate the amount of information contained within a graph xu2023better . We compute the properties of one graph and concatenate them to get a graph-level representation.
Self-supervised annotator. To assign labels to graphs in a self-supervised manner, we employ clustering algorithms on the graph-level representations. The number of clusters is determined jointly by the silhouette score rousseeuw1987silhouettes and the separation of the graphs. The candidates of the number of clusters are chosen to ensure different clusters are well separated. Among the candidates, we select the final number of clusters by maximizing the mean Silhouette Coefficient of all samples.
Next we detail the parameterization of the denoising model $\tilde{p}_{\theta}(\widetilde{\mathbf{A}}^{0}\mid\mathbf{A}^{t})$ with the self-assigned graph-level labels $\mathbf{k}$. The denoising model recovers the edges of the original adjacency matrix by predicting the connectivity of the upper triangle, which can be formulated as a link prediction problem zhang2018link ; kumar2020link . Following the link prediction setup, the denoising model is composed of a graph transformer (GT) shi2020masked and an MLP link predictor. Denote the hidden dimension as $d$, we treat the node degrees as node features and utilize a linear mapping $f_{d}:\mathbb{R}\mapsto\mathbb{R}^{d}$ to match the dimension. Similarly, we utilize another linear mapping $f_{t}:\mathbb{R}\mapsto\mathbb{R}^{d}$ for timestep $t$ and learnable embeddings $f_{k}:\{0,\dots,K\}\mapsto\mathbb{R}^{d}$ for labels $\mathbf{k}$, where $K$ is the number of clusters. The outputs are summed together and then fed into the GT. Mathematically, we have
$\begin{split}&\mathbf{h}^{t}=\operatorname{GT}\left(f_{d}\left(\operatorname{degree}\left(\mathbf{A}^{t}\right)\right)+f_{t}(t)+f_{k}(\mathbf{k}),\mathbf{A}^{t}\right),\\ &\tilde{p}_{\theta}(\widetilde{\mathbf{A}}^{0}_{ij}\mid\mathbf{A}^{t};t,\mathbf{k})\coloneq\tilde{p}_{\theta}(\widetilde{\mathbf{A}}^{0}_{ij}\mid\mathbf{h}^{t})=\operatorname{MLP}\left([\mathbf{h}^{t}_{i},\mathbf{h}^{t}_{j}]\right).\end{split}$ (4)
With the above denoising network, our diffusion model is trained on the collected graphs by optimizing the variational lower bound in (3). After the pre-training process, we perform adaptive downstream enhancement through graph structure augmentation.
3.3 Downstream adaptation through data augmentation
The downstream phase of UniAug is to augment the graph topology through guided generation. This guidance process serves to provide downstream semantics for the diffusion model, thus bridging the gap between the pre-training distribution and the downstream datasets. Among the techniques for diffusion guidance, gradient-based methods dhariwal2021diffusion ; gruver2024protein offer versatile approaches by incorporating external conditions that are not present during training. For the discrete diffusion process, we opt for the gradient-based NOS method gruver2024protein due to its flexibility and efficiency. Specifically, we build an MLP regression head $g_{\theta}:\mathbb{R}^{d}\mapsto\mathbb{R}^{r}$ that takes the hidden representations $\mathbf{h}^{t}$ as the input and outputs the guidance objective of dimension $r$. Denote $\tau$ as the temperature, $\gamma$ as the step-size, $\lambda$ as the regularization strength, and $\varepsilon$ drawn from $\mathcal{N}(0,I)$, we sample from $\tilde{p}^{\prime}(\widetilde{\mathbf{A}}^{0}\mid\mathbf{h}^{t})\propto\tilde{p}_{\theta}(\widetilde{\mathbf{A}}^{0}\mid\mathbf{h}^{t})\exp\left(g_{\theta}\left(\mathbf{h}^{t}\right)\right)$ via Langevin dynamics
$\mathbf{h}^{t,\prime}\leftarrow\mathbf{h}^{t,\prime}-\gamma\nabla_{\mathbf{h}^{t,\prime}}\left[\lambda\mathrm{KL}\left(\tilde{p}^{\prime}\left(\widetilde{\mathbf{A}}^{0}\mid\mathbf{h}^{t,\prime}\right)\|\ \tilde{p}^{\prime}\left(\widetilde{\mathbf{A}}^{0}\mid\mathbf{h}^{t}\right)\right)-g_{\theta}\left(\mathbf{h}^{t,\prime}\right)\right]+\sqrt{2\gamma\tau}\varepsilon.$ (5)
One key question to answer is how to choose the proper guidance objectives. Our goal is to find numerical characteristics that can best describe the structural properties of a graph. This includes supervision signal and self-supervised information on the level of node, edge, and graph.
Node level. Node labels provide the supervision signal for node classification tasks. Beyond node labels, node degrees are a fundamental factor in the evolutionary process of a graph liu2011controllability . From the perspective of network analysis, centrality measures indicate the importance of nodes from various viewpoints borgatti2005centrality . Empirically, we observe that utilizing different node-level heuristics as guidance targets tends to yield similar outcomes. Therefore, we focus on node labels and node degrees.
Edge level. Edge-level heuristics can be broadly classified into two categories: local structural heuristics, such as Common Neighbor and Adamic Adar adamic2003friends , and global structural heuristics, such as Katz katz1953new and SimRank jeh2002simrank . Similar to node-level heuristics, empirical observations suggest that different edge-level heuristics tend to yield comparable guidance effects. In this work, we focus on the Common Neighbors (CN) heuristic due to its efficiency. Another edge-level guidance objective is to recover the adjacency matrix from the node representations in a link prediction way, similar to how we parameterize the denoising network. We anticipate that such link prediction objective helps to align the generated graph with the downstream data on the granularity of edges.
Graph level. Graph labels offer the supervision signal for graph classes or regression targets. In addition, we incorporate graph-level properties xu2023better as quantitative measures to bridge the gap between the pre-training distribution and the downstream dataset. We empirically observe that graph label guidance offers significantly higher performance boosts compared to properties on graph-level tasks. Therefore, we focus on graph labels in our experiments.
We provide our choice of objectives for each task in Appendix B. We note that all the above objectives are natural choices inspired by heuristics and downstream tasks. There exist many other self-supervised objectives to be explored, such as community-level spectral change tan2024community and motif occurrence prediction rong2020self . We leave the study of these objectives as one future work. With the diffusion guidance, we assemble the augmented graphs with generated structures and original node features. The augmented graphs are then fed into downstream-specific GNNs.
3.4 Comparison to existing methods
Table 1: Comparison between GDA methods, pre-training methods, and UniAug. By cross-domain transfer, we emphasize the ability of the method to train on vastly different domains and benefit all of them.
GDA methods Pre-training methods UniAug
GraphAug CFLP Half-Hop FLAG AttrMask D-SLA GraphMAE
Effective on graph-level task
Effective on edge-level task
Effective on node-level task
In-domain transfer
Cross-domain transfer
The data augmentation paradigm of UniAug allows us to disentangle the upstream and downstream. We construct a diffusion model as the upstream component to comprehend the structural patterns of graphs across various domains. In addition, we leverage downstream inductive biases with downstream-specific models in a plug-and-play manner. This allows UniAug to facilitate cross-domain transfer, offering a unified method that benefits graphs across different domains for various downstream tasks. On the contrary, existing GDA methods are typically designed for specific tasks and hard to transfer to unseen patterns. In the meantime, existing pre-training methods fail to transfer across domains due to heterogeneity in features and structures. This comparison highlights the success of UniAug as a data-scaling graph structure augmentor across domains. We summarize the comparison between methods in Table 1.
4 Experiment
In this section, we conduct experiments to validate the effectiveness of UniAug. We first pre-train our discrete diffusion model on thousands of graphs collected from diverse domains. For each downstream task, we train an MLP guidance head with corresponding objectives on top of the diffusion model. We then perform structure augmentation using UniAug and subsequently train a task-specific GNN on augmented data for prediction. Through the experiments, we aim to answer the following research questions:
  • RQ1: Can UniAug benefit graphs from various domains across different downstream tasks?
  • RQ2: What is the scaling behavior of UniAug corresponding to data scale and amount of compute?
  • RQ3: Which components of UniAug are effective in preventing negative transfer?
4.1 Main results
To get a comprehensive understanding of UniAug, we evaluate it on $25$ downstream datasets from $7$ domains for graph property prediction, link prediction, and node classification. The statistics of the datasets can be found in Appendix C, and technical details of the experiments are in Appendix D.
Baselines. We evaluate our model against three main groups of baselines. (1) Task-specific GNNs: For graph property prediction, we use GIN xu2018powerful ; for link prediction, we use GCN kipf2017semisupervised and NCN wang2024neural ; and for node classification, we use GCN kipf2017semisupervised . (2) Graph pre-training methods: These include AttrMask, CtxtPred, EdgePred, and InfoMax hu2019strategies , JOAO you2021graph , D-SLA kim2022graph , and GraphMAE hou2022graphmae . For each of these methods, we pre-train it on the same pre-training set as UniAug. While most of the pre-training graphs lack node features, we calculate the node degrees as the input. Each method consists of three pre-trained variants with different backbone GNNs, including GIN, GCN, and GAT. We note that all these methods require the downstream graphs to have the same node feature space as the pre-training data. Therefore, in the fine-tuning stage, we replace the node features of the downstream datasets with node degrees, evaluate all three variants, and report the highest performance for each method in each task. (3) Graph data augmentation (GDA) methods: For graph property prediction, we include S-Mixup ling2023graph , GraphAug luo2022automated , FLAG kong2022robust , GREA liu2022graph , and DCT liu2024data ; for link prediction, we include CFLP zhao2022learning ; and for node classification on heterophilic graphs, we include Half-Hop azabou2023half . The GDA methods are implemented based on chosen task-specific GNNs.
Table 2: Mean and standard deviation of accuracy (%) with $10$-fold cross-validation on graph classification. The best result is bold. The highlighted results indicate negative transfer for pre-training methods compared to GIN. The last column is the average rank of each method.
DD Enzymes Proteins NCI1 IMDB-B IMDB-M Reddit-B Reddit-12K Collab A.R.
GIN 75.81 ± 6.11 66.00 ± 7.52 73.32 ± 4.03 78.30 ± 3.20 71.10 ± 2.90 49.07 ± 2.81 90.85 ± 1.30 48.63 ± 1.62 74.54 ± 2.41 5.56
AttrMask 72.93 ± 3.09 23.66 ± 6.09 73.10 ± 3.90 77.67 ± 2.53 71.20 ± 2.40 48.00 ± 3.14 87.50 ± 3.31 48.00 ± 1.60 75.64 ± 1.52 8.00
CtxtPred 75.14 ± 2.67 21.67 ± 3.87 72.21 ± 4.60 78.99 ± 1.29 70.70 ± 1.55 48.20 ± 2.23 90.35 ± 2.31 47.62 ± 2.50 75.60 ± 1.49 7.67
EdgePred 75.64 ± 2.77 22.00 ± 3.32 71.22 ± 3.53 77.82 ± 2.95 70.20 ± 2.23 47.80 ± 2.42 90.80 ± 1.69 48.35 ± 1.44 74.64 ± 2.24 8.56
InfoMax 75.23 ± 3.43 22.50 ± 6.76 71.30 ± 5.18 76.94 ± 1.48 71.60 ± 2.06 46.70 ± 2.46 89.15 ± 2.84 48.98 ± 1.83 75.44 ± 1.12 8.00
JOAO 75.98 ± 2.86 22.17 ± 3.67 71.57 ± 5.31 76.87 ± 2.27 71.02 ± 1.81 48.85 ± 2.06 90.17 ± 2.13 49.01 ± 1.90 74.77 ± 1.71 7.11
D-SLA 74.66 ± 3.30 22.67 ± 4.21 71.97 ± 4.17 77.95 ± 2.11 71.92 ± 2.75 47.28 ± 1.88 89.77 ± 1.87 48.50 ± 1.33 75.99 ± 2.08 7.00
GraphMAE 76.07 ± 3.25 23.00 ± 3.64 70.45 ± 4.19 79.08 ± 2.72 71.50 ± 2.01 47.93 ± 3.03 86.10 ± 3.63 47.67 ± 1.16 74.84 ± 1.36 7.67
S-Mixup 73.12 ± 3.27 66.85 ± 7.04 74.61 ± 5.08 78.91 ± 1.61 69.61 ± 4.43 48.33 ± 5.36 88.65 ± 3.12 48.30 ± 2.50 75.89 ± 3.26 6.67
GraphAug 75.21 ± 2.63 68.14 ± 7.92 74.21 ± 3.70 79.53 ± 3.21 74.00 ± 3.41 48.11 ± 1.85 90.50 ± 3.17 49.00 ± 1.99 76.02 ± 2.67 3.67
FLAG 76.87 ± 7.21 68.35 ± 7.45 74.31 ± 4.21 79.03 ± 3.75 68.83 ± 4.67 47.21 ± 3.45 89.11 ± 2.40 47.48 ± 3.01 75.32 ± 3.13 7.00
UniAug 78.13 ± 2.61 71.50 ± 5.85 75.47 ± 2.50 80.54 ± 1.77 73.50 ± 2.48 50.13 ± 2.05 92.28 ± 1.59 49.48 ± 0.71 77.00 ± 2.02 1.11
Table 3: Mean and standard deviation of MAE $\downarrow$ across $10$ runs on molecule regression. The last column is the average rank of each method. Among the methods, all pre-training methods discard atom and bond features due to dimension mismatch and we include the best-performing method JOAO into comparison; GIN and UniAug remove the bond features; others incorporate both.
ogbg-Lipo ogbg-ESOL ogbg-FreeSolv A.R.
GINE* 0.545 ± 0.019 0.766 ± 0.016 1.639 ± 0.146 5.00
GIN 0.543 ± 0.021 0.729 ± 0.018 1.613 ± 0.155 3.67
JOAO 0.859 ± 0.007 1.458 ± 0.040 3.292 ± 0.117 7.00
FLAG* 0.528 ± 0.012 0.755 ± 0.039 1.565 ± 0.098 3.00
GREA* 0.586 ± 0.036 0.805 ± 0.135 1.829 ± 0.368 6.00
DCT* 0.516 ± 0.071 0.717 ± 0.020 1.339 ± 0.075 1.33
UniAug 0.528 ± 0.006 0.677 ± 0.026 1.448 ± 0.049 1.67
Graph property prediction. We employ graph label guidance for UniAug throughout the graph-level tasks by training a $2$-layer MLP as the guidance head on the graph labels in the training set. In the augmentation stage, we generate multiple graphs per training sample, and the generated graphs are then fed into the baseline GIN. We present the results of molecule regression in Table 3 and graph classification in Table 2. Three key observations emerge from the analysis: (1) Existing pre-training methods show negative transfer compared to GIN. Some special cases are the Enzymes and molecule regression datasets, where all pre-training methods fail to yield satisfactory results. In these datasets, the features are one of the driving components for graph property prediction, while the pre-training methods fail to encode such information due to incompatibility with the feature dimension. This reveals one critical drawback of the pre-training methods: their inability to handle feature heterogeneity. (2) GDA methods yield inconsistent results across different datasets. While these methods enhance performance in some datasets, they cause performance declines in others. This variability is directly reflected in the average rank, where some of them even fall behind the GIN. (3) Unlike the pre-training methods and GDA methods, UniAug shows consistent performance improvements against GIN with a large margin. In the molecule regression tasks, UniAug effectively compensates for the absence of bond features and achieves performance comparable to DCT, which is a data augmentation method pre-trained on in-domain molecule graphs. These findings affirm that the pre-training and structure augmentation paradigm of UniAug effectively benefits the downstream datasets at the graph level.
Link prediction. We choose three guidance objectives for UniAug, including node degree, CN, and link prediction objective, as described in Section 3.3. For each objective, we train an MLP to provide guidance information. We then augment the graph structure by generating a synthetic graph and preserving the original training edges, ensuring that the augmented graph does not remove any existing edges. The augmented graph is then fed into a GCN for link prediction. We summarize the results in Table 4, which show similar patterns to those observed in graph property prediction: (1) Existing pre-training methods provide negative transfer, especially on datasets with node features. (2) GDA method CFLP leads to performance drops on the datasets without features and also suffers from high computation complexity during preprocessing. (3) UniAug enhances performance across all tested datasets. In addition, we employ UniAug to NCN wang2024neural , one of the state-of-the-art methods for link prediction. The results demonstrate consistent performance boosts from UniAug when we apply NCN as the backbone. The structure augmentation paradigm of UniAug allows plug-and-play applications to any downstream-specific models, showcasing its adaptability and effectiveness. In addition, we study the effects of three guidance objectives. More details can be found in Appendix D.2.
Table 4: Mean and standard deviation across $10$ runs on link prediction. Results are scaled ×$\times 100$. The last two methods are based on NCN, while the rest are GCN-based. The best result is bold for two backbones, respectively. The highlighted results indicate negative transfer for pre-training methods compared to GCN. The last column is the average rank of each GCN-based method.
Cora Citeseer Pubmed Power Yeast Erdos Flickr A.R.
MRR MRR MRR Hits@10 Hits@10 Hits@10 Hits@10
GCN 30.26 ± 4.80 50.57 ± 7.91 16.38 ± 1.30 30.61 ± 4.07 24.71 ± 4.92 35.71 ± 2.65 8.10 ± 2.58 4.14
AttrMask 13.43 ± 1.93 20.23 ± 1.29 16.39 ± 3.62 29.92 ± 2.61 25.10 ± 4.77 30.85 ± 3.13 8.77 ± 1.65 6.43
CtxtPred 15.68 ± 2.91 22.31 ± 1.31 13.10 ± 3.70 29.30 ± 3.55 22.96 ± 4.28 34.82 ± 2.55 3.61 ± 1.01 7.86
EdgePred 15.31 ± 3.54 22.91 ± 1.87 17.85 ± 4.45 29.54 ± 3.78 25.78 ± 4.51 34.65 ± 3.84 6.86 ± 3.24 5.43
InfoMax 16.35 ± 2.57 22.90 ± 1.30 15.91 ± 2.71 29.29 ± 4.72 26.33 ± 4.12 35.82 ± 4.12 3.23 ± 0.38 6.00
JOAO 17.21 ± 3.66 23.10 ± 1.41 15.33 ± 3.70 28.98 ± 4.01 26.47 ± 4.65 33.77 ± 3.05 6.01 ± 1.57 6.00
D-SLA 15.55 ± 3.12 23.05 ± 1.54 16.10 ± 3.96 29.37 ± 2.88 26.15 ± 3.32 36.02 ± 4.58 6.70 ± 2.03 5.29
GraphMAE 15.94 ± 1.73 20.35 ± 1.52 13.80 ± 1.36 27.69 ± 1.99 26.51 ± 2.92 35.63 ± 3.61 8.41 ± 2.44 6.14
CFLP 33.62 ± 6.44 55.20 ± 4.16 17.01 ± 2.75 16.02 ± 8.31 24.23 ± 5.23 28.74 ± 2.38 OOM 6.43
UniAug-GCN 35.36 ± 7.88 54.66 ± 4.55 17.28 ± 1.89 34.36 ± 1.68 27.52 ± 4.80 39.67 ± 4.51 9.46 ± 1.18 1.29
NCN 31.72 ± 4.48 58.03 ± 3.45 38.26 ± 2.56 27.36 ± 5.00 39.85 ± 5.07 36.81 ± 3.29 8.33 ± 0.92
UniAug-NCN 35.92 ± 7.85 61.69 ± 3.21 40.30 ± 2.53 30.20 ± 1.46 42.11 ± 5.74 39.26 ± 2.84 8.85 ± 0.90
Table 5: Mean and standard deviation of accuracy (%) across $10$ splits on node classification of heterophilic graphs. The best result is bold. The highlighted results indicate negative transfer for pre-training methods compared to GCN. The last column is the average rank of each method.
Cornell Wisconsin Texas Actor Chameleon* Squirrel* A.R.
GCN 59.41 ± 6.03 51.68 ± 4.34 63.78 ± 4.80 30.58 ± 1.29 40.94 ± 3.91 39.11 ± 1.74 3.83
AttrMask 44.86 ± 5.43 53.73 ± 4.31 60.54 ± 5.82 25.31 ± 1.03 35.81 ± 2.88 30.63 ± 1.68 5.83
CtxtPred 40.81 ± 7.78 36.67 ± 17.23 58.92 ± 4.32 23.97 ± 2.63 24.36 ± 4.13 26.26 ± 7.50 9.50
EdgePred 42.70 ± 5.51 48.04 ± 6.63 59.37 ± 5.11 22.99 ± 6.22 21.02 ± 5.06 27.94 ± 8.41 8.83
InfoMax 39.19 ± 12.75 39.80 ± 16.38 58.87 ± 4.06 23.30 ± 4.37 22.59 ± 4.91 27.52 ± 9.09 10.17
JOAO 40.13 ± 8.60 44.70 ± 7.45 57.06 ± 3.43 24.17 ± 5.02 25.81 ± 3.79 31.72 ± 7.03 8.33
D-SLA 41.05 ± 6.88 42.13 ± 9.58 59.93 ± 4.29 23.74 ± 4.06 26.49 ± 4.27 28.50 ± 6.90 8.00
GraphMAE 47.05 ± 4.37 57.06 ± 4.59 63.70 ± 5.51 24.69 ± 0.68 37.18 ± 3.08 31.94 ± 1.65 5.00
Half-Hop 62.46 ± 7.58 76.47 ± 2.61 72.35 ± 4.27 33.95 ± 0.68 38.59 ± 2.89 37.34 ± 2.18 3.00
UniAug 68.11 ± 6.72 69.02 ± 4.96 73.51 ± 5.06 33.11 ± 1.57 43.84 ± 3.39 41.90 ± 1.90 2.00
UniAug + Half-Hop 72.43 ± 5.81 79.61 ± 5.56 77.03 ± 4.27 34.97 ± 0.55 41.94 ± 2.77 38.79 ± 2.61 1.50
  • *Chameleon and Squirrel are filtered to remove duplicated nodes platonov2023a .
Table 6: Results of node classification on homophily graphs. Results are scaled ×$\times 100$.
Cora Citeseer Pubmed
ACC $\uparrow$ GCN 81.75 ± 0.73 70.71 ± 0.76 79.53 ± 0.25
UniAug 81.78 ± 0.60 71.17 ± 0.58 79.54 ± 0.35
SD $\downarrow$ GCN 24.51 ± 1.06 22.57 ± 0.80 27.02 ± 0.56
UniAug 23.45 ± 0.90 19.90 ± 0.81 26.50 ± 0.55
Node classification. To demonstrate the effectiveness of UniAug in node-level tasks, we transform the node classification into subgraph classification. Specifically, we extract the aggregation tree of each node, i.e., $2$-hop subgraph for a $2$-layer GCN, and label the subgraph with the center node. We then adopt a strategy similar to graph classification and train a $2$-layer classifier as a guidance head. Inspired by the success of structure augmentation on heterophilic graphs bi2022make ; azabou2023half , we evaluate UniAug on $6$ heterophilic datasets. We observe phenomena similar to those seen in graph property prediction and link prediction in Table 5. One thing to mention is the combination of UniAug and Half-Hop. Half-Hop offers performance improvements in four out of six datasets via data augmentation, and combining it with UniAug yields even higher results. This highlights the flexibility of UniAug and opens up possibilities for further exploration of its use cases. Given the impressive results of UniAug on heterophilic graphs, we anticipate it will also help to balance the performance disparities among nodes with different homophily ratios on homophilic graphs mao2024demystifying . We split the nodes into five groups according to their homophily ratios and calculate the standard deviation (SD) across groups. As shown in Table 6, UniAug matches the performance of vanilla GCN and also reduces the performance discrepancies corresponding to SD.
Figure 3: Effects of pre-training data scale on graph classification (left) and link prediction (right). The groups SMA, FUL, and EXT represent SMALL, FULL, and EXTRA data collection.
[…]
4.2 Scaling behavior of UniAug
In light of the neural scaling law kaplan2020scaling ; hoffmann2022training ; abnar2022exploring ; zhai2022scaling ; liu2024neural , we expect UniAug to benefit from an increased volume of data and more compute budget. In this subsection, we investigate the scaling behavior of UniAug in terms of data scale and amount of compute for pre-training.
Data scale During the data collection process, we prepare three versions of the training data with increasing magnitude. We first sample $10$ graphs per category from the Network Repository rossi2015network to build a SMALL collection. Next, we gather all the graphs from the Network Repository and filter out large-scale graphs and outliers for a FULL collection. In addition, we add a $1000$ graphs subset of the GitHub Star dataset from TUDataset morris2020tudataset to enlarge the coverage of diverse patterns and form an EXTRA collection. We pre-train three versions of UniAug respectively on the three collections and evaluate them on graph classification and link prediction. As shown in Fig. 3, we observe a clear trend of increase in performance as we enlarge the scale of pre-training data. This paves the way to scale up UniAug to even more pre-training graphs.
Amount of compute We sought to understand how effectively our diffusion model can learn data patterns as we continue to train it. To this end, we checkpointed UniAug every 2,000 epochs ($5\times 10^{-3}$ PF-days) while training on the EXTRA collection, and then applied it to graph classification and link prediction tasks. The results are illustrated in Fig. 4. We observe that downstream performance generally improves with prolonged training, while the trend slows down for some datasets when we reach 8,000 epochs. We take the checkpoint at the 10,000th epoch for evaluations. Given the scaling behavior observed, we anticipate UniAug to become even more effective with additional resources.
4.3 Preventing negative transfer
Table 7: Demonstration of negative transfer on graph classification (up) and link prediction (down).
Enzymes Proteins IMDB-B IMDB-M
GIN 66.00 ± 7.52 73.32 ± 4.03 71.10 ± 2.90 49.07 ± 2.81
UniAug 71.50 ± 5.85 75.47 ± 2.50 73.50 ± 2.48 50.13 ± 2.05
w/o self-cond 71.11 ± 7.50 73.31 ± 4.63 71.50 ± 2.27 49.00 ± 2.74
w/o guidance 62.17 ± 3.93 71.15 ± 4.56 53.80 ± 3.29 35.33 ± 3.17
w/ cross-guide 51.50 ± 7.64 72.46 ± 4.35 71.10 ± 2.38 49.20 ± 2.59
Cora Citeseer Power Yeast Erdos
MRR MRR Hits@10 Hits@10 Hits@10
GCN 30.26 ± 4.80 50.57 ± 7.91 30.61 ± 4.07 24.71 ± 4.92 35.71 ± 2.65
UniAug 35.36 ± 7.88 54.66 ± 4.55 34.36 ± 1.68 27.52 ± 4.80 39.67 ± 4.51
w/o self-cond 27.97 ± 16.11 37.65 ± 6.00 28.95 ± 7.73 23.54 ± 8.28 34.33 ± 6.18
w/o guidance 29.60 ± 6.06 51.41 ± 7.10 25.57 ± 6.04 25.26 ± 6.06 37.11 ± 4.16
w/ cross-guide 32.37 ± 4.20 50.59 ± 5.67 32.99 ± 2.54 26.76 ± 3.88 36.30 ± 3.67
In the previous parts of the experiments, we showcase the positive transfer of UniAug across different tasks. We now investigate which aspects of the design prevent negative transfer. UniAug consists of two main components: a pre-trained diffusion model and the structure augmentation through guided generation. In the pre-training process, we inject self-supervised graph labels into the diffusion model and we wonder about the performance of its unconditioned counterpart. Regarding the augmentation process, we examine the impact of diffusion guidance by exploring outcomes when the guidance is either removed or applied using another dataset from a different domain (cross-guide). We summarize the results in Table 7 for graph classification and link prediction. All modifications investigated lead to performance declines in both tasks. We observe that removing guidance results in significant negative transfers for graph classification, while the effects of self-conditioning are more pronounced for link prediction. We conclude that both the self-conditioning strategy and diffusion guidance are crucial in preventing negative transfer, underscoring their importance in the design of UniAug.
5 Conclusion and Discussion
In this work, we propose a graph structure augmentation pipeline UniAug to leverage the increasing scale of graph data. We collect thousands of graphs from various domains and pre-train a self-conditioned discrete diffusion model on them. In the downstream stage, we augment the graphs by preserving the original node features and generating synthetic structures. We apply UniAug to node-, link-, and graph-level tasks and achieve consistent performance gain. We have successfully developed a showcase that benefits from cross-domain graph data scaling using diffusion models.
One limitation of the current analysis is the absence of an investigation into the effects of model parameters due to limited resources. Given the scaling behavior of UniAug in terms of data scale and amount of compute, we anticipate that a large-scale model will provide significant performance improvements. One future direction is to investigate the adaptation of fast sampling methods to the discrete diffusion models on graphs. This will lead to lower time complexity and enable broader application scenarios.
Appendix A Derivation of Diffusion Process
In the following, we will formulate the existing discrete diffusion models into binary diffusion on the adjacency matrix. We denote the adjacency matrix of a graph as $\mathbf{A}^{0}\in\{0,1\}^{n\times n}$ with $n$ nodes. Following D3PM [49], we corrupt the adjacency matrix into a sequence of latent variables $\mathbf{A}^{1:T}=\mathbf{A}^{1},\mathbf{A}^{2},\dots,\mathbf{A}^{T}$ by independently injecting noise into each element with a Markov process
$q\left(\mathbf{A}^{t}\mid\mathbf{A}^{t-1}\right)=\prod_{i,j:i<j}\operatorname{Cat}\left(\mathbf{A}^{t}_{ij};\mathbf{p}=\mathbf{A}^{t-1}_{ij}\mathbf{Q}^{t}\right),$ (6)
where $\mathbf{Q}^{t}\in[0,1]^{2\times 2}$ is the transition probability of timestep $t$. The above Markov process is called forward process. Existing works provide different designs for the transition matrix $\mathbf{Q}^{t}$, including
[…]
where $\pi$ is the converging non-zero probability and $\beta^{t}$ is the noise scale. All three transition matrices can be written as binary diffusion with Bernoulli distribution
$\begin{split}q\left(\mathbf{A}^{t}\mid\mathbf{A}^{t-1}\right)&=\operatorname{Bernoulli}\left(\mathbf{A}^{t};\alpha^{t}\mathbf{A}^{t-1}+\left(1-\alpha^{t}\right)\pi\right),\\ q\left(\mathbf{A}^{t}\mid\mathbf{A}^{0}\right)&=\operatorname{Bernoulli}\left(\mathbf{A}^{t};\bar{\alpha}^{t}\mathbf{A}^{0}+\left(1-\bar{\alpha}^{t}\right)\pi\right),\\ q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{t},\mathbf{A}^{0}\right)&=\frac{q\left(\mathbf{A}^{t}\mid\mathbf{A}^{t-1}\right)q\left(\mathbf{A}^{t-1}\mid\mathbf{A}^{0}\right)}{q\left(\mathbf{A}^{t}\mid\mathbf{A}^{0}\right)},\end{split}$ (8)
where $\alpha^{t}=1-\beta^{t}$ and $\bar{\alpha}^{t}=\prod_{i=1}^{t}\alpha^{i}$. The prior $\mathbf{A}^{T}$ is determined by $\pi$ with $p\left(\mathbf{A}^{T}_{ij}\right)=\operatorname{Bernoulli}(\pi)$, i.e., the existence of each edge follows a Bernoulli distribution with probability $\pi$. The main difference of the forward process among the existing works is the choice of $\pi$, where $\pi=0$ for EDGE [47], $\pi=0.5$ for D4Explainer [48], and a pre-computed average density $\pi$ for DiGress [46].
In our early experiments, we observe that the absorbing kernel $\pi=0$ surpasses the other two in terms of efficiency and effectiveness for graph generation. The forward process with non-zero $\pi$ will add non-existing edges, which brings in additional computations. When sampling from prior, non-zero $\pi$ will introduce additional uncertainty because we will first sample every edge from $\operatorname{Bernoulli}(\pi)$. Therefore, we choose the absorbing prior $\pi=0$ in this work and leave the exploration of other transition kernels as a future work.
We note that in our implementation, we choose the number of timesteps $T$ as $128$ according to our early experiments and some existing works [77, 47]. We leave the study of the effects of diffusion timesteps on downstream tasks as a future work.
Appendix B Guidance Objective for Downstream Tasks
We mention various guidance objectives in Section 3.3 with different granularity. Here, we specify the objectives we use for each downstream task. Our empirical results suggest that supervision signals will lead to better performance. Thus, we use node labels for node classification and graph labels for graph property prediction in Section 4. Regarding link prediction, we anticipate that both node-level and edge-level objectives may help the downstream adaptation. Therefore, we choose three objectives including node degree, CN heuristic, and link prediction objective.
Appendix C Datasets
The license of the datasets use in this work is in Table 8.
Table 8: List of datasets and corresponding License
Dataset License
Network Repository CC BY-SA
Github Star CC BY 4.0
Cora NLM license
Citeseer NLM license
Pubmed NLM license
WebKB MIT license
Wikipedia Network MIT license
Actor MIT license
Power BSD License
Yeast BSD License
Erdos BSD License
Amazon Photo MIT license
Flickr MIT license
DD CC BY 4.0
Enzymes CC BY 4.0
Proteins CC BY 4.0
NCI1 CC BY 4.0
IMDB CC BY 4.0
Reddit CC BY 4.0
Graph property prediction datasets include DD and Proteins [78], Enzymes [79], NCI1 [80], IMDB-Binary, IMDB-Multi, Reddit-Binary, and Reddit-Multi-12K [81], ogbg-Lipo, ogbg-ESOL and ogbg-FreeSolv [9]. The statistics are summarized in 9.
Table 9: Statistics of graph property prediction datasets.
Domain Dataset Task type # Graphs # Tasks # Nodes # Edges
Biology DD Classification 1,178 2 284 716
Enzymes Classification 600 6 33 64
Proteins Classification 1,113 2 40 73
Academic Collab Classification 5,000 3 74 2458
Social IMDB-B Classification 1,000 2 20 97
IMDB-M Classification 1,500 3 13 66
Reddit-5k Classification 4,999 5 509 595
Reddit-12k Classification 11,929 11 391 1305
Chemical NCI Classification 4,110 2 30 32
ogbg-Lipo Regression 4200 1 27 59
ogbg-ESOL Regression 1128 1 13 27
ogbg-FreeSolv Regression 642 1 9 17
Link prediction datasets include Cora, Citeseer, and Pubmed [82], Power [83], Yeast [84], Erdos [85], Amazon Photo [86], and Flickr [10]. The statistics are summarized in 10.
Table 10: Statistics of link prediction datasets.
Cora Citeseer Pubmed Power YST ERD Flickr
Domain Citation Transport Biology Academic Social
#Nodes 2,708 3,327 18,717 4,941 2,284 6,927 334,863
#Edges 5,278 4,676 44,327 6,594 6,646 11,850 899,756
Mean Degree 3.9 2.81 4.74 2.67 5.82 3.42 5.69
Node classification datasets include Cora, Citeseer, and Pubmed [82], WebKB (Texas, Cornell, and Wisconsin) [87], Wikipedia Network (Chameleon and Squirrel) [87], and Actor [88]. The first three are homophilic graphs, and the others are heterophilic. The statistics are summarized in 11.
Table 11: Statistics of node classification datasets.
Cora Citeseer Pubmed Cornell Wisconsin Texas Chameleon* Squirrel* Actor
Domain Citation Web Social
#Nodes 2,708 3,327 19,717 183 251 183 890 2,223 7,600
#Edges 5,278 4,676 44,324 295 499 309 8,854 46,998 33,544
#Classes 7 6 3 5 5 5 5 5 5
  • *Chameleon and Squirrel are filtered to remove duplicated nodes [73].
Appendix D Experiment
In this section, we introduce the implementation details and additional results for the experiments. Throughout all the experiments, we train all the methods with Adam optimizer on an A100 GPU. We train the guidance head of UniAug with cross-entropy loss for class labels and mean squared error loss for all other objectives. For multi-class objectives, we apply the label smoothing [89] technique following NOS [21]. Denote $y$ as the one-hot label and $C$ as the number of classes, we have
$\overline{\mathbf{y}}_{t}=\bar{\alpha}_{t}*\mathbf{y}+\left(1-\bar{\alpha}_{t}\right)/C*\mathbf{1}.$ (9)
D.1 Graph property prediction
For graph classification, we follow [90] for the setting with $10$-fold cross-validation. We utilize a $5$-layer GIN with latent dimensions of $64$ throughout the datasets. For molecule regression, we implement a $5$-layer GIN with a virtual node, and the latent dimensions are $300$. We have mainly four hyperparameters for UniAug: step-size $\gamma$ and regularization strength $\lambda$ in (5), number of repeats per training graph, and whether augment validation and test graphs with the trained guidance head. For each training graph, we repeatedly generate structures and plug in the original node features for multi-repeat augmentation. We perform the update in (5) for $5$ times per each sampling step. The hyperparameters are tuned from the choices in Table 12.
Table 12: Hyperparameter choices for graph property prediction.
$\lambda$ 0.01
$\gamma$ [0.1, 0.5, 1.0]
# repeats [1, 5, 10, 32, 64]
Aug val and test [True, False]
D.2 Link prediction
For link prediction, we follow the model designs and evaluation protocols of [91]. For results based on GCN and NCN, we use a GCN encoder to produce node embeddings and perform link prediction with a prediction head. The prediction head of GCN is a $3$-layer MLP. The number of layers and the latent dimension of the GCN encoder are taken from [91]. We have mainly three hyperparameters for UniAug: step-size $\gamma$ and regularization strength $\lambda$, and the number of updates in (5) per each sampling step. In addition, inspired by the pseudo labeling strategy [92], we provide an option threshold $q$ for the sampling process of the diffusion model. Specifically, we only keep the edges with the probability of existence higher than $q$ for each sampling step. After the sampling process, we recover the training edges of the original graph structure. The hyperparameters are tuned from the choices in Table 13.
Table 13: Hyperparameter choices for link prediction.
$\lambda$ [0.01, 1, 100]
$\gamma$ [0.1, 1.0, 10.0]
$q$ [None, 0.9, 0.99, 0.999]
# updates [5, 10, 20]
One thing to mention is that we handle the large graphs by graph partitioning with METIS [93]. Specifically, we augment the partitions of a large graph and then assemble the partitions back into a single graph. The edges between different partitions are recovered after the assembling process.
Table 14: Effects of different guidance objectives.
Cora Citeseer Pubmed Power Yeast Erdos Flickr
MRR MRR MRR Hits@10 Hits@10 Hits@10 Hits@10
Link guide 30.45 ± 2.90 54.66 ± 4.55 16.97 ± 0.92 33.41 ± 2.95 25.80 ± 4.10 36.79 ± 1.98 9.46 ± 1.18
Degree guide 32.73 ± 6.71 51.13 ± 5.51 16.37 ± 0.58 32.88 ± 2.02 27.52 ± 4.80 39.67 ± 4.51 9.11 ± 0.88
CN guide 35.36 ± 7.88 50.86 ± 5.73 17.28 ± 1.89 34.36 ± 1.68 26.67 ± 4.02 36.18 ± 4.32 9.28 ± 1.18
As mentioned in Section 4.1, we choose three guidance objectives for link prediction with different granularity. The effects of different objectives can be found in Table 14. We observe that the outcomes of different objectives differ across datasets and there is no consistently winning strategy.
D.3 Node classification
For node classification on heterophilic graphs, we use the fixed splits from Geom-GCN [87] for Cornell, Wisconsin, Texas, and Actor. For Chameleon and Squirrel, we remove duplicated nodes following [73] and take their fixed splits. Regarding node classification on homophilic graphs, we employ the semi-supervised setting [94]. The GCN backbone is implemented as a $2$-layer classifier. Similar to graph property prediction, we have mainly four hyperparameters for UniAug: step-size $\gamma$ and regularization strength $\lambda$ in (5), number of repeats per training graph, and whether augment validation and test graphs with the trained guidance head. The hyperparameters are tuned from the choices in Table 15.
Table 15: Hyperparameter choices for node classification.
$\lambda$ 0.01
$\gamma$ [0.1, 0.5, 1.0]
# repeats [1, 5, 10]
Aug val and test [True, False]
D.4 Investigation on scaling
In Section 4.2, we investigate the scaling behavior of UniAug regarding data scale and pre-training time. We omit some of the results for a better visualization. Here we present the numerical results in Table 16 and Table 17.
Table 16: Effects of pre-training data scale on graph classification (up) and link prediction (down).
Enzymes Proteins IMDB-B IMDB-M
GIN 66.00 ± 7.52 73.32 ± 4.03 71.10 ± 2.90 49.07 ± 2.81
UniAug- SMALL 66.83 ± 7.38 73.50 ± 5.61 69.80 ± 2.70 48.93 ± 3.20
UniAug- FULL 71.33 ± 6.51 74.05 ± 4.82 73.11 ± 2.35 49.67 ± 2.41
UniAug- EXTRA 71.17 ± 7.10 75.47 ± 2.50 73.50 ± 2.48 50.13 ± 2.05
Cora Citeseer Power Yeast Erdos
MRR MRR Hits@10 Hits@10 Hits@10
GCN 30.26 ± 4.80 50.57 ± 7.91 30.61 ± 4.07 24.71 ± 4.92 35.71 ± 2.65
UniAug- SMALL 32.25 ± 8.71 47.91 ± 3.87 32.25 ± 3.72 25.81 ± 4.89 36.28 ± 3.56
UniAug- FULL 32.81 ± 7.44 48.32 ± 6.00 32.97 ± 3.75 26.36 ± 4.62 36.07 ± 4.20
UniAug- EXTRA 35.36 ± 7.88 54.66 ± 4.55 34.36 ± 1.68 27.52 ± 4.80 39.67 ± 4.51
Table 17: Effects of pre-training amount of compute on graph classification (up) and link prediction (down).
$10^{-3}$ PF-days Enzymes Proteins IMDB-B IMDB-M
5 68.18 ± 6.21 73.32 ± 3.63 71.20 ± 2.90 48.28 ± 2.75
10 69.00 ± 5.10 74.30 ± 5.33 72.80 ± 3.85 48.60 ± 2.23
15 68.83 ± 5.88 75.11 ± 3.18 71.77 ± 2.38 48.60 ± 2.48
20 70.79 ± 5.73 74.87 ± 5.30 73.04 ± 2.82 49.47 ± 2.20
25 71.50 ± 5.85 75.47 ± 2.50 73.50 ± 2.48 50.13 ± 2.05
$10^{-3}$ PF-days Cora Citeseer Power Yeast Erdos
MRR MRR Hits@10 Hits@10 Hits@10
5 27.56 ± 4.36 49.45 ± 9.20 22.81 ± 9.47 23.62 ± 9.77 35.33 ± 3.16
10 31.02 ± 6.53 50.72 ± 6.22 32.49 ± 2.52 26.70 ± 4.85 36.10 ± 4.66
15 33.24 ± 7.97 49.02 ± 5.92 32.88 ± 3.31 27.80 ± 4.55 39.70 ± 3.67
20 34.71 ± 9.08 52.90 ± 3.84 33.69 ± 3.23 26.90 ± 3.93 39.33 ± 3.16
25 35.36 ± 7.88 54.66 ± 4.55 34.36 ± 1.68 27.52 ± 4.80 39.67 ± 4.51
Appendix E Broader Impact
In this work, we build a universal graph structure augmentor that benefits from data scaling across domains. Given the consistent performance improvements for different tasks, we expect this work to contribute significantly towards the goal of building a graph foundation model. In the meantime, we showcase the power of the deep generative models on graphs by introducing new application scenarios. We anticipate such success will contribute to the community of generative models and graph learning.
It is important to mention that the model backbones of our method and baselines heavily rely on neighboring node information as an inductive bias. However, this characteristic can result in biased predictions, especially when patterns in neighborhood majorities dominate, leading to potential ethical issues in model predictions.